Allison disposable income is $90,000, her autonomous consumption is $15,000 and her Marginal Propensity to Consume is 0.80, how much money is Allison saving?

Select one:

a. She is saving $3,000

b. She is saving $7,000

c. She is saving $10,000

d. She is saving $12,000

e. She is saving $15,000
Answer #1

Allison's disposable income= $90000

Allison's autonomous consumption = $ 15000

Income available for her induced consumption = $90000 - $15000 = $75000

Allison's MPC = 0.80

Therefore her MPS = 1 - 0.80 = 0.20 ( As MPC+MPS=1)

allioson's saving = $75000 * 0.20 = $15000